OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6.

OpenStudy (tkhunny):

Have you considered the Distributive Property to eliminate the parentheses?

OpenStudy (anonymous):

3x +3= -2x+2+6 3x+3+-2x+8 3x+2x=8-3 5x=5 x=1